# Multilingual Evaluation of Long Context Retrieval and Reasoning

### v1

@@ Abstract
Recent large language models (LLMs) demonstrate impressive capabilities in handling long contexts, some exhibiting near-perfect recall on synthetic retrieval tasks. However, these evaluations have mainly focused on English text and involved a single target sentence within lengthy contexts. Our work investigates how LLM performance generalizes to multilingual settings with multiple hidden target sentences. We comprehensively evaluate several long-context LLMs on retrieval and reasoning tasks across five languages: English, Vietnamese, Indonesian, Swahili, and Somali. These languages share the Latin script but belong to distinct language families and resource levels. Our analysis reveals a significant performance gap between languages. The best-performing models such as Gemini-1.5 and GPT-4o, achieve around 96% accuracy in English to around 36% in Somali with a single target sentence. However, this accuracy drops to 40% in English and 0% in Somali when dealing with three target sentences. Our findings highlight the challenges long-context LLMs face when processing longer contexts, an increase in the number of target sentences, or languages of lower resource levels.
Multilingual Evaluation of Long Context Retrieval and Reasoning
@@ 1 Introduction
Long-context models, particularly multilingual ones, have the potential to enable remarkable progress in various applications by understanding lengthy textual data across different languages. An example of this potential was recently demonstrated by the newly introduced Gemini-1.5 Pro model (Reid et al., 2024) which leveraged its long-context window for in-context learning. By including a grammar manual in its context window, the model was able to learn to translate from English to Kalamang, an extremely low-resource language with fewer than 200 speakers (Visser, 2020). Such examples highlight the potential of long-context models in tackling challenging tasks in low-resource languages, where data scarcity has traditionally been a barrier.
@@
To address this gap, we present the first comprehensive study of long-context LLMs in multilingual settings leveraging evaluation frameworks relying on synthetic tasks (Mohtashami and Jaggi, 2023; Chen et al., 2023; Liu et al., 2024; Kamradt, 2023; Reid et al., 2024; Anthropic, 2024). Although the task is partially synthetic, we create a new dataset [^1] The dataset will be made available for facilitating further research. consisting of naturally occurring text and human translated data, making the setup as close to real-world setting while creating a controlled environment for comparing model performance across languages. In addition to retrieval tasks, we introduce a new reasoning task where the models not only need to retrieve relevant items but also compare them with each other. For this, the models must keep track of these items in a long-context scenario, allowing us to analyze the models’ reasoning capabilities.
We conduct a systematic evaluation of six different LLMs across five languages of varying resource levels. Our research aims to answer the following two questions: (1) How do the long context capabilities of LLMs compare in retrieval and reasoning tasks in multilingual contexts? (2) Are there significant performance differences between LLMs in multilingual contexts?
@@ 2 Related Work
Although some long-context real-world benchmarks have been recently introduced, they are also limited to English (An et al., 2023), and while some bilingual English/Chinese (Bai et al., 2023; Qiu et al., 2024; Yuan et al., 2024) datasets offer a slight improvement, due to the effort-intensive nature of dataset creation, they are limited to a very small number of languages.
@@ 3.1 Languages and Models
Language ISO 639-3 Code Resource Level Language Family Script
English eng Level 5 Indo-European Latin
Vietnamese vie Level 4 Austro-Asiatic Latin
Indonesian ind Level 3 Austronesian Latin
Swahili swa Level 2 Niger-Congo Latin
Somali som Level 1 Afro-Asiatic Latin
Table 1: Languages studied and their details
Languages  We selected five languages to study: English, Vietnamese, Indonesian, Swahili, and Somali. These languages span different resource levels from high to extremely low [^2] The linguistic diversity taxonomy (Joshi et al., 2020) is used to identify the resource levels. allowing us to gain insight into how language resource levels affect models’ ability to work over long context windows. To deliberately control for script-related variables, we chose languages that use the Latin script for three reasons: models perform significantly better on Latin-script languages than non-Latin languages (Chau and Smith, 2021; Bang et al., 2023), the fragmentation rate of Latin script is lower than other scripts allowing Latin-script languages to be represented with substantially fewer tokens as compared to languages in other scripts (Ács, 2019; Ahia et al., 2023) – a disparity that becomes even more pronounced over long contexts, and the fragmentation rate of Latin-script languages remains comparable which is helpful when considering long input texts. Our selection of languages, shown in Table 1, has the added benefit of including less-studied languages and language families [^3] Language families were obtained from Ethnologue: Languages of the World, available at https://www.ethnologue.com/., providing a more comprehensive view of the latest generation of multilingual capabilities of long-context LLMs.
@@ 3.2 Retrieval and Reasoning Tasks
Language models with the ability to handle long context rely heavily on their capacity to retrieve relevant information from the given text and reason based on that information to interpret and follow human instructions effectively. Although synthetic tasks alone may not provide a comprehensive assessment of a language model’s long-context capabilities, they offer the advantage of being easily adaptable to specific scenarios and languages. This is particularly important given that the most recent long-context real-world benchmarks are limited to English (An et al., 2023) or bilingual English/Chinese (Bai et al., 2023; Qiu et al., 2024; Yuan et al., 2024). Moreover, there is some evidence to suggest that performance on synthetic retrieval tasks can, to a certain extent, generalize to real-world datasets (Qiu et al., 2024). Therefore, carefully designed synthetic tasks can serve as a valuable tool for evaluating a language model’s long-context capabilities across a diverse range of languages.
@@ 3.2.1 Retrieving a Needle ($n=1$)
In this task, the model’s objective is to locate and extract information from a single target sentence hidden somewhere in the haystack. We adopt the same needle pattern as used in previous studies (Dhinakaran, 2024; Reid et al., 2024; Anthropic, 2024), which takes the form: “The special magic {city} number is: {number}”. Here, {city} is randomly chosen from a list of 69 unique cities from around the world, and {number} is a randomly generated 7-digit number. The list of cities were automatically translated and then post-edited into all the languages and are available in Appendix A.
@@ 3.2.2 Reasoning over Multiple Needles ($n>1$)
Figure 1: Ablation results of comparing Paul Graham’s essays and news articles serving as haystacks for English experiments tested using GPT-4 model
3.3 Dataset: mLongRR
Previous work has extensively relied on Paul Graham’s essays in English to make up the haystacks. Translating these essays into multiple languages could potentially introduce translation errors. Instead, we create a new dataset – mLongRR – by collecting BBC news articles published in the five languages [^6] Our language selection was primarily constrained by the availability of authentic texts, especially those within the same script but varying in resource levels.. This approach allows us to work with sufficiently long, naturally occurring, with the added benefit that this recently published data is less likely to have been encountered by the models during training. To assess the impacts of different data for haystacks, we conducted an ablation study using the GPT-4 model and observed no noticeable differences in the model performance when the haystacks consisted of Paul Graham’s essays or news articles in English, as shown in Figure 1.
@@ 3.4 Prompts
In our initial run of experiments, we explored two existing prompt templates used in previous work: prompt 1 (Kamradt, 2023; Dhinakaran, 2024; Anthropic, 2024) and prompt 2 (Reid et al., 2024).
Figure 2: Ablation results of comparing two different prompts
@@
Although the articles in mLongRR are already in the native languages, we hired four professional human translators from the freelancing website fiverr.com to translate the needles, city names, and prompts from English into Vietnamese, Indonesian, Swahili, and Somali. In our preliminary experiments with the GPT-4 model, we experimented with English and language-specific prompts, and found that English prompts worked better than language-specific prompts, in line with previous studies (Etxaniz et al., 2023; Barreiß et al., 2024; Lai et al., 2023b), for three of four non-English languages (with the exception of Swahili). As a result, the remaining experiments were conducted using English prompts.
@@ 3.6 Evaluation
The responses generated by the models were used to calculate the accuracy. For both the retrieval and reasoning tasks, the models generated a short, straightforward text containing the 7-digit number (for number-based reasoning) or the city name (for city-based reasoning). For example, a typical output looked like this: "3210496" or "The larger magic number is 8134445". We extracted the number/city and compared it to the ground truth to check whether the model’s response was correct or not. For languages other than English, the models occasionally responded with the city name in English or the target language and both were acceptable.
@@ 4.1 Performance of different models across languages and tasks
Average 1.946 1.538 2.24 2.346 1.776 1.484
Table 2: Tokenization rate for each language using different model tokenizers
4.2 Performance across varying needle depths and haystack lengths
Figure 5: Language-specific heatmap visualizations with varying depths on the $y$-axis and context lengths on the $x$-axis, averaged over all the models, when $n=1$, $n=2$, and $n=3$
@@ 4.3 Performance across different languages
The results presented in Figure 5 provide a fine-grained analysis of language-specific performance. English consistently performs well across tasks, with near 100% accuracy in simpler tasks but declining with increased complexity, particularly at greater depths and longer contexts. Vietnamese also maintains high accuracy, though it declines similarly with complexity. Indonesian starts reasonably well but drops significantly in more complex scenarios. Swahili shows weaker overall performance, struggling with all tasks, especially complex ones. Somali performs the poorest, often reaching zero accuracy as task complexity increases, highlighting challenges in handling this low-resource language. In short, performance degrades progressively as we move from high-resource languages to low-resource languages. The detailed results of each model and language are included in Appendix B.
@@ 4.4 Impact of tokenization
We further analyze the tokenization rate (the average number of tokens generated per word in each language) of the different languages, as presented in Table 2. Unsurprisingly, English consistently shows the lowest tokenization rates across all models. Vietnamese has varying rates, with Gemini-1.5 being the most efficient, while Claude-3 and YaRN-7b tokenize more heavily. Indonesian exhibits moderate rates with some variability across models. Swahili and Somali, the two lowest resource level languages in our study, generally have higher tokenization rates, suggesting these are more challenging for models to process effectively.
@@ 4.5 Reasoning about magic numbers and world cities
Lastly, we compare the performance of the models in reasoning tasks with 2 and 3 needles for two types of question prompts: identifying the larger/largest magic number (e.g., 4281932) or the city with the larger/largest magic number (e.g., Doha). From the results presented in Figure 6, we observe that the models yield generally better performance in the “number” tasks compared to the “city” tasks implying that they may be more adept at handling numerical reasoning than reasoning over geographic entities. This suggests that improvements in model training might be necessary to enhance performance in tasks involving non-numerical data, particularly in diverse and low-resource languages.
Figure 6: Comparing reasoning over magic numbers and random cities, when $n=2$ and $n=3$
@@ 5 Conclusion
Appendix A List of cities used in needles
The following lists the cities [^7] https://github.com/Arize-ai/LLMTest_NeedleInAHaystack/blob/main/LLMNeedleHaystackSynthesis.py#L341 used to generate the random needles along with translations in Vietnamese, Indonesian, Swahili, and Somali.
English Names: Almaty, Amman, Amsterdam, Antananarivo, Astana, Athens, Baghdad, Bangalore, Bangkok, Barcelona, Beirut, Belgrade, Berlin, Bratislava, Brussels, Bucharest, Budapest, Buenos Aires, Cairo, Chicago, Colombo, Copenhagen, Dakar, Damascus, Delhi, Doha, Dubai, Helsinki, Ho Chi Minh City, Istanbul, Jakarta, Johannesburg, Kampala, Khartoum, Kigali, Kuala Lumpur, Lagos, Lima, Lisbon, Los Angeles, Madrid, Maputo, Melbourne, Mexico City, Moscow, Mumbai, Nairobi, Oslo, Paris, Port Louis, Rabat, San Francisco, Santiago, Sarajevo, Seattle, Seoul, Shanghai, Sofia, Sydney, Tashkent, Thimphu, Tokyo, Toronto, Tunis, Vancouver, Victoria, Vienna, Yangon, Yerevan
Vietnamese Names: Almaty, Amman, Amsterdam, Antananarivo, Astana, Athens, Baghdad, Bangalore, Bangkok, Barcelona, Beirut, Belgrade, Berlin, Bratislava, Brussels, Bucharest, Budapest, Buenos Aires, Cairo, Chicago, Colombo, Copenhagen, Dakar, Damascus, Delhi, Doha, Dubai, Helsinki, Istanbul, Jakarta, Johannesburg, Kampala, Khartoum, Kigali, Kuala Lumpur, Lagos, Lima, Lisbon, Los Angeles, Madrid, Maputo, Melbourne, Moskva, Mumbai, Nairobi, Oslo, Paris, Port Louis, Rabat, San Francisco, Santiago, Sarajevo, Seattle, Seoul, Sofia, Sydney, Tashkent, Thimphu, Thành phố Hồ Chí Minh, Thành phố México, Thượng Hải, Tokyo, Toronto, Tunis, Vancouver, Victoria, Vienna, Yangon, Yerevan
Indonesian Names: Almaty, Amman, Amsterdam, Antananarivo, Astana, Athena, Baghdad, Bangkok, Barcelona, Beirut, Bengaluru, Beograd, Berlin, Bratislava, Brussel, Budapest, Buenos Aires, Bukares, Chicago, Dakar, Damaskus, Delhi, Doha, Dubai, Helsinki, Ho Chi Minh, Istanbul, Jakarta, Johannesburg, Kairo, Kampala, Khartoum, Kigali, Kolombo, Kopenhagen, Kopenhagen , Kota Meksiko, Kuala Lumpur, Lagos, Lima, Lisboa, Los Angeles, Madrid, Maputo, Melbourne, Moskwa, Mumbai, Nairobi, Oslo, Paris, Port Louis, Rabat, San Francisco, Santiago, Sarajevo, Seattle, Seoul, Shanghai, Sofia, Sydney, Tashkent, Thimphu, Tokyo, Toronto, Tunis, Vancouver, Victoria, Wina, Yangon, Yerevan
Swahili Names: Almaty, Amman, Amsterdam, Antananarivo, Astana, Athene, Baghdad, Bangkok, Barcelona, Beirut, Belgrade, Bengaluru, Berlin, Bratislava, Brussels, Budapest, Buenos Aires, Bukarest, Chicago, Colombo, Dakar, Damasko, Delhi, Doha, Dubai, Helsinki, Ho Chi Minh, Istanbul, Jakarta, Jiji la Meksiko, Johannesburg, Kairo, Kampala, Khartoum, Kigali, Kopenhagen, Kuala Lumpur, Lagos, Lima, Lisbon, Los Angeles, Madrid, Maputo, Melbourne, Moskow, Mumbai, Nairobi, Oslo, Paris, Port Louis, Rabat, San Francisco, Santiago, Sarajevo, Seattle, Seoul, Shanghai, Sofia, Sydney, Tashkent, Thimphu, Tokyo, Toronto, Tunis, Vancouver, Victoria, Vienna, Yangon, Yerevan
Somali Names: Almaty, Amman, Amsterdam, Antananarivo, Astana, Athens, Bangkok, Baqdad, Barcelona, Beirut, Belgrade, Bengaluru, Berlin, Bratislava, Brussels, Bucharest, Budapest, Buenos Aires, Chicago, Colombo, Copenhagen, Dakar, Delhi, Dimishiq, Doha, Dubai, Helsinki, Ho Chi Minh, Istanbul, Jakarta, Johannesburg, Kampala, Khartuum, Kigali, Kuala Lumpur, Lagos, Lima, Lisbon, Los Angeles, Madrid, Magaalada Mexico, Maputo, Melbourne, Moscow, Mumbai, Nairobi, Oslo, Paris, Port Louis, Qaahira, Rabat, San Francisco, Santiago, Sarajevo, Seattle, Seoul, Shanghai, Sofia, Sydney, Tashkent, Thimphu, Tokyo, Toronto, Tunis, Vancouver, Victoria, Vienna, Yangon, Yerevan
Appendix B Detailed model and language-specific results

### v3

# Evaluating Multilingual Long-Context Models for Retrieval and Reasoning

@@ Abstract
Recent large language models (LLMs) demonstrate impressive capabilities in handling long contexts, some exhibiting near-perfect recall on synthetic retrieval tasks. However, these evaluations have mainly focused on English text and involved a single target sentence within lengthy contexts. Our work investigates how LLM performance generalizes to multilingual settings with multiple hidden target sentences. We create a new dataset – mLongRR – to comprehensively evaluate several multilingual long-context LLMs on retrieval and reasoning tasks across five languages: English, Vietnamese, Indonesian, Swahili, and Somali. These languages share the Latin script but belong to distinct language families and resource levels. Our analysis reveals a significant performance gap between languages. The best-performing models such as Gemini-1.5 and GPT-4o, achieve around 96% accuracy in English to around 36% in Somali with a single target sentence. However, this accuracy drops to 40% in English and 0% in Somali when dealing with three target sentences. Our findings highlight the challenges long-context LLMs face when processing longer contexts, an increase in the number of target sentences, or languages of lower resource levels.
Evaluating Multilingual Long-Context Models for Retrieval and Reasoning
@@ 1 Introduction
Long-context language models, particularly multilingual ones, have the potential to enable remarkable progress in various applications by understanding lengthy textual data across different languages. An example of this potential was recently demonstrated by the newly introduced Gemini-1.5 Pro model (Reid et al., 2024) which leveraged its long-context window for in-context learning. By including a grammar manual in its context window, the model was able to learn to translate from English to Kalamang, an extremely low-resource language with fewer than 200 speakers (Visser, 2020). Such examples highlight the potential of long-context models in tackling challenging tasks in low-resource languages, where data scarcity has traditionally been a barrier.
@@
Language ISO 639-3 Code Resource Level Language Family Script
English eng Level 5 Indo-European Latin
Vietnamese vie Level 4 Austro-Asiatic Latin
Indonesian ind Level 3 Austronesian Latin
Swahili swa Level 2 Niger-Congo Latin
Somali som Level 1 Afro-Asiatic Latin
Table 1: Languages studied and their details.
To address this gap, we present the first comprehensive study of long-context LLMs in multilingual settings leveraging evaluation frameworks relying on synthetic tasks (Mohtashami and Jaggi, 2023; Chen et al., 2023; Liu et al., 2024; Kamradt, 2023; Reid et al., 2024; Anthropic, 2024). Although the task is partially synthetic, we create a new dataset – mLongRR [^1] The code is available at https://github.com/PortNLP/mLongRR. – consisting of naturally occurring text and human translated data, making the setup as close to real-world setting while creating a controlled environment for comparing model performance across languages. In addition to retrieval tasks, we introduce a new reasoning task where the models not only need to retrieve relevant items but also compare them with each other. For this, the models must keep track of these items in a long-context scenario, allowing us to analyze the models’ reasoning capabilities.
We conduct a systematic evaluation of six different LLMs across five languages of varying resource levels. Our research aims to answer the following two questions:
(1) How do the long context capabilities of LLMs compare in retrieval and reasoning tasks in multilingual contexts?
(2) Are there significant performance differences between LLMs in multilingual contexts?
@@ 2 Related Work
Although some long-context real-world benchmarks have been recently introduced, they are also limited to English (An et al., 2024), and while some bilingual English/Chinese (Bai et al., 2023; Qiu et al., 2024; Yuan et al., 2024) datasets offer a slight improvement, due to the effort-intensive nature of dataset creation, they are limited to a very small number of languages.
@@ 3.1 Languages and Models
Languages  We selected five languages to study: English, Vietnamese, Indonesian, Swahili, and Somali. These languages span different resource levels from high to extremely low [^2] The linguistic diversity taxonomy (Joshi et al., 2020) is used to identify the resource levels. allowing us to gain insight into how language resource levels affect models’ ability to work over long context windows.
Furthermore, we deliberately control for script-related variables as they have been shown to have a considerable impact on the performance of a model Bagheri Nezhad and Agrawal (2024b). We study languages that use the Latin script for three reasons: models perform significantly better on Latin-script languages than non-Latin languages (Chau and Smith, 2021; Bang et al., 2023), the fragmentation rate of Latin script is lower than other scripts allowing Latin-script languages to be represented with substantially fewer tokens as compared to languages in other scripts (Ács, 2019; Ahia et al., 2023) – a disparity that becomes even more pronounced over long contexts, and, lastly, the fragmentation rate of Latin-script languages remains comparable which is helpful when considering considerably long input texts. Our selection of languages, shown in Table 1, has the added benefit of including less-studied languages and language families [^3] Language families were obtained from Ethnologue: Languages of the World, available at https://www.ethnologue.com/., providing a more comprehensive view of the latest generation of multilingual capabilities of long-context LLMs.
@@ 3.2 Retrieval and Reasoning Tasks
Language models with the ability to handle long context rely heavily on their capacity to retrieve relevant information from the given text and reason based on that information to interpret and follow human instructions effectively. Although synthetic tasks alone may not provide a comprehensive assessment of a language model’s long-context capabilities, they offer the advantage of being easily adaptable to specific scenarios and languages. This is particularly important given that the most recent long-context real-world benchmarks are limited to English (An et al., 2024) or bilingual English/Chinese (Bai et al., 2023; Qiu et al., 2024; Yuan et al., 2024). Moreover, there is some evidence to suggest that performance on synthetic retrieval tasks can, to a certain extent, generalize to real-world datasets (Qiu et al., 2024). Therefore, carefully designed synthetic tasks can serve as a valuable tool for evaluating a language model’s long-context capabilities across a diverse range of languages.
@@ 3.2.1 Retrieving a Needle ($n=1$)
In this task, the model’s objective is to locate and extract information from a single target sentence hidden somewhere in the haystack. We adopt the same needle pattern as used in previous studies (Dhinakaran, 2024; Reid et al., 2024; Anthropic, 2024), which takes the form: “The special magic {city} number is: {number}”. Here, {city} is randomly chosen from a list of 69 unique cities from around the world, and {number} is a randomly generated 7-digit number. The list of cities were automatically translated and then post-edited into all the languages.
@@ 3.2.2 Reasoning over Multiple Needles ($n>1$)
Figure 1: Ablation results of comparing Paul Graham’s essays and news articles serving as haystacks for English experiments tested using GPT-4 model.
3.3 Creating mLongRR Dataset
Prior work has extensively relied on Paul Graham’s essays in English to make up the haystacks (Kamradt, 2023; Dhinakaran, 2024; Anthropic, 2024). Translating these essays into multiple languages, however, could potentially introduce translation errors. Instead, we create a new dataset – mLongRR – by collecting BBC news articles published in the five languages [^6] Our language selection was primarily constrained by the availability of authentic texts, especially those within the same script but varying in resource levels., inspired by recent work Bagheri Nezhad and Agrawal (2024a). This approach allows us to work with sufficiently long, naturally occurring, with the added benefit that this recently published data is less likely to have been encountered by the models during training. It is worth noting that this does not result in a parallel dataset as the news articles are often specific to their respective regions.
We assess the impacts of different data for haystacks by conducting an ablation study using the GPT-4 model. As shown in Figure 1, we observed no noticeable differences in the model performance when the haystacks consisted of Paul Graham’s essays or news articles in English. To statistically confirm this observation, we applied McNemar’s test McNemar (1947), which yielded a p-value of 1.0, indicating no significant difference between the two datasets.
@@ 3.4 Prompts
In our initial run of experiments, we explored two existing prompt templates used in previous work:
  • prompt 1 (Kamradt, 2023; Dhinakaran, 2024; Anthropic, 2024), and
  • prompt 2 (Reid et al., 2024).
Figure 2: Ablation results of comparing two different prompts.
@@
Although the articles in mLongRR are already in the native languages, we hired four professional human translators from the freelancing website fiverr.com to translate the needles, city names, and prompts from English into Vietnamese, Indonesian, Swahili, and Somali. In our preliminary experiments with the GPT-4 model, we experimented with English and language-specific prompts, and found that English prompts worked better than language-specific prompts, in line with previous studies (Etxaniz et al., 2024; Bareiß et al., 2024; Lai et al., 2023b), for three of four non-English languages (with the exception of Swahili). As a result, the remaining experiments were conducted using English prompts.
@@ 3.6 Evaluation
The responses generated by the models were used to calculate the accuracy. For both the retrieval and reasoning tasks, the models generated a short, straightforward text containing the 7-digit number (for number-based reasoning) Dhinakaran (2024) or the city name (for city-based reasoning). For example, a typical output looked like this: "3210496" or "The larger magic number is 8134445". We extracted the number/city and compared it to the ground truth to check whether the model’s response was correct or not. For languages other than English, the models occasionally responded with the city name in English or the target language and both were acceptable.
@@ 4.1 Performance of different models across languages and tasks
Figure 5: Language-specific heatmap visualizations with varying depths on the $y$-axis and context lengths on the $x$-axis, averaged over all the models, when $n=1$, $n=2$, and $n=3$.
4.2 Performance across varying needle depths and haystack lengths
@@
Average 1.94 1.53 2.24 2.34 1.77 1.48
Table 2: Tokenization rate for each language using different model tokenizers.
@@ 4.3 Performance across different languages
The results presented in Figure 5 provide a fine-grained analysis of language-specific performance. English consistently performs well across tasks, with near 100% accuracy in simpler tasks but declining with increased complexity, particularly at greater depths and longer contexts. Vietnamese also maintains high accuracy, though it declines similarly with complexity. Indonesian starts reasonably well but drops significantly in more complex scenarios. Swahili shows weaker overall performance, struggling with all tasks, especially complex ones. Somali performs the poorest, often reaching zero accuracy as task complexity increases, highlighting challenges in handling this low-resource language. In short, performance degrades progressively as we move from high-resource languages to low-resource languages. The detailed results of each model and language are included in Appendix A.
@@
Figure 6: Comparing reasoning over magic numbers and random cities, when $n=2$ and $n=3$ (results obtained using Gemini-1.5).
@@ 4.4 Impact of tokenization
We further analyze the tokenization rate, also known as fertility rate, which is the average number of tokens generated per word for the different languages. The results are presented in Table 2. Unsurprisingly, English consistently shows the lowest tokenization rates across all models. Vietnamese has varying rates, with Gemini-1.5 being the most efficient, while Claude-3 and YaRN-7b tokenize more heavily. Indonesian exhibits moderate rates with some variability across models. Swahili and Somali, the two lowest resource level languages in our study, generally have higher tokenization rates, suggesting these are more challenging for models to process effectively.
@@ 4.5 Reasoning about magic numbers and world cities
Lastly, we compare the performance of the models in reasoning tasks with 2 and 3 needles for two types of question prompts: identifying the larger/largest magic number (e.g., 4281932) or the city with the larger/largest magic number (e.g., Doha). From the results presented in Figure 6, we observe that the models yield generally better performance in the “number” tasks compared to the “city” tasks implying that they may be more adept at handling numerical reasoning than reasoning over geographic entities, however, this trend is reversed for Swahili and Somali.
@@ 5 Conclusion
Acknowledgments
We thank the anonymous reviewers as well as the members of PortNLP lab for their insightful comments that helped improve this paper. This research was partially supported by the National Science Foundation grant HNDS-R 2242205.
Appendix A Detailed model and language-specific results